# Alpha-element abundance patterns in star-forming regions of the local Universe

### v1

@@ 5.3 Ar/O
Figure 13: log(Ar/O) as a function of 12+log(O/H) (top) and as a function of 12+log(Ar/H) (bottom). Black squares and blue circles represent star-forming galaxies and H ii regions of the DESIRED-E sample, which Ar/O ratios have been calculated using the ICF(Ar) scheme by Izotov et al. (2006). The orange diamonds indicate the mean log(Ar/O) values considering bins in 12+log(O/H) or 12+log(Ar/H). The red continuous lines represent linear fits to the data represented in each panel. The green continuous line shows the time evolution of Ar and O abundances of the ISM predicted by the CEM of the Milky Way of Medina-Amayo (2023) and Amayo & Carigi (in preparation). The black dashed lines and the grey bands show the solar 12+log(O/H), 12+log(Ar/H) and log(Ar/O) and their associated uncertainties, respectively, from Asplund et al. (2021).
In Sect. 4.3, we concluded that the different ICF(Ar) schemes considered show a very similar behaviour of the Ar/O ratio with respect to the O abundance and degree of ionisation, so the analysis of the Ar/O ratio seems to be independent of the final selection of the ICF(Ar) scheme (see the small differences in the Ar/H computed with the different schemes in Fig 14). We finally decided to use the ICF(Ar) by Izotov et al. (2006) in order to be consistent with the results obtained with the rest of alpha-elements and to eliminate the objects with a lower degree of ionisation (O2+/O $<$ 0.2), due to the larger contribution of unseen Ar+ that could increase the dispersion of those data points.
@@
In Fig. 13 we represent the log(Ar/O) vs. 12+log(O/H) (top) and vs. 12+log(Ar/H) (bottom) distributions of SFGs+H ii regions of the DESIRED-E sample. We have represented the data points with the same symbol regardless of whether [Ar iii] or [Ar iii]$+$[Ar iv] lines were used to determine the Ar abundance. As in some previous figures, in Fig. 13 we also include a least-squares fit to the data represented in each panel. The two distributions show weak correlations with slopes of different signs. Although the sample Pearson correlation coefficients of the linear fits to log(Ar/O) vs. 12+log(O/H) and vs. 12+log(Ar/H) distributions are r = --0.28 and 0.17, respectively, their very low $p$-values indicate that both correlations can be considered real. As in the case of the fits separated by type of object, the fit to log(Ar/O) vs. 12+log(O/H) for SFGs+H ii regions indicates a decrease of log(Ar/O) of 0.24 dex in the whole interval of 12+log(O/H) represented in the top panel of Fig. 13. This behaviour is contradictory to what is expected from current nucleosynthesis and chemical evolution models. In fact, the curve predicted by the CEM of Medina-Amayo (2023) and Amayo & Carigi (in preparation) included in Fig. 13 would imply an increase of the Ar/O ratio with metallicity, in clear disagreement with the general behavior shown by our data. This puzzling result deserves further study
@@
In the case of the log(Ar/O) vs. 12+log(Ar/H) distribution shown in Fig. 13, we find a weak linear correlation with a small positive slope and a total increase of log(Ar/O) of 0.12 dex in the entire Ar/H range. Although the predictions of the CEM are now qualitatively consistent with the data, they are not quantitatively consistent. The CEM predicts an increase of 0.54 dex over the range of Ar abundances represented in Fig. 13, about four time higher than observed. On the other hand, unlike what happens in the cases of Ne and S, the CEM also does not reproduce the mean log(Ar/O) at low metallicities. The predicted value is about 0.10 dex lower than observed, but this cannot possibly be due to depletion of O onto dust grains because it is expected to be negligible at such low metallicities.
@@
Regarding the log(Ar/O) vs. 12+log(Ar/H) distribution obtained by other authors, the results are disparate. Arellano-Córdova et al. (2024) and Miranda-Pérez & Hidalgo-Gámez (2023) obtain a quite flat distribution in the log(Ar/O) vs. 12+log(Ar/H) relation at intermediate values of 12+log(Ar/H) but opposite trends in low and high values of the Ar abundance. In particular, Miranda-Pérez & Hidalgo-Gámez (2023) find a rather strong increase of log(Ar/O) for objects with 12+log(Ar/H) $>$ 6.0, reaching values larger than in any SFG of our DESIRED-E sample. On the other hand, Arnaboldi et al. (2022) studied the behaviour of the Ar/O vs. Ar/H relation for a sample of planetary nebulae in the M31 galaxy. They find an increase of the Ar/O ratio as the Ar/H increases but in a rather narrow interval of 12+log(Ar/H), between 6.0 and 6.5 approximately, in agreement with the predictions of the chemical evolution models of the Milky Way thick disc and solar neighbourhood of Kobayashi et al. (2020). Arnaboldi et al. (2022) interpret the increase in the log(Ar/O) vs. 12+log(Ar/H) plane they observe as due to the contribution of SNe Ia. In fact, the models by Kobayashi et al. (2020) predict that about 34% of the total solar Ar is produced by SNe Ia. Fig. 16 also indicates that in the CEM of Medina-Amayo (2023) and Amayo & Carigi (in preparation) the increase of Ar/O with metallicity is produced basically by SNe Ia. We do not see an Ar/O increase as large as that reported by Arnaboldi et al. (2022), and our results are broadly consistent with a constant log(Ar/O) distribution in most of the 12+log(Ar/H) range. However, the bins represented in the bottom panel of our Fig. 13 suggest a slight increase –of only $\sim$0.1 dex– for values of 12+log(Ar/H) $\gtrsim$ 6.0 which –as we demonstrate in Fig. 12– does not appear to be an artifact of the ICF(Ar). In principle, our linear fit to the log(Ar/O) vs. 12+log(Ar/H) distribution could be interpreted as an indication of a certain contribution by SNe Ia, although in a more limited proportion than that proposed by Arnaboldi et al. (2022).
@@ 6 Conclusions
We derive Ne/O ratios for 1228 DESIRED-E spectra, of those 557 correspond to H ii regions and 671 to SFGs. The log(Ne/O) vs. 12+log(O/H) distribution of H ii regions shows a large dispersion and no evidence of correlation. Our analysis indicates that the position of those objects cannot be correctly reproduced by the ICF(Ne) schemes considered. On the other hand, the SFGs show much lower dispersion in their Ne/O ratios and rather similar log(Ne/O) vs. 12+log(O/H) and log(Ne/O) vs. 12+log(Ne/H) linear relations with positive but small slopes. These correlations may be due to several factors or a combination of them. Firstly, the effect of a metallicity-dependent O depletion fraction that increases throughout the metallicity range covered by the objects. Secondly, an effect of the ICF(Ne), which does not completely cancel out the Ne/O dependence on O2+/O. A third possibility would be a certain unknown mechanism producing Ne, but chemical evolution models do not predict such scenario, at least in a significant amount.
The log(S/O) vs. 12+log(O/H) relationship that we found separately for the 441 and 492 spectra of H ii regions and SFG of our sample is compatible with a constant value, especially for H ii regions and when we combine all the sample objects (H ii regions + SFGs), in contrast with some previous works (Dors et al., 2016; Díaz & Zamora, 2022; Brazzini et al., 2024). On this regard, we have found that at least part of the results of Díaz & Zamora (2022) are incorrect, so some of their conclusions should be revised. We find a rather tight and positive linear fit for all the objects when we use the log(S/O) vs. 12+log(S/H) distribution. It is remarkable that the predictions of the CEM by Medina-Amayo (2023) and Amayo & Carigi (in preparation) fits very well that distribution in the whole S/H range. We propose that the positive slope the log(S/O) vs. 12+log(S/H) relation and its change of slope at 12+log(S/H) $\sim$ 6.0 can be interpreted as an observational evidence of contribution of S produced by SNe Ia.
@@ Appendix B The relation between the oxygen abundance and ionisation degree
Figure 15: 12+log(O/H) as a function of the ionisation degree, O2+/O, for DESIRED-E sample of Galactic and extragalactic H ii regions (blue circles, top panel) and star-forming galaxies (black squares, bottom panel). The red diamonds and lines indicate the mean 12+log(O/H) values and its standard deviation, considering bins of 0.2 in O2+/O.
Appendix C Relative contribution to O, Ne, S and Ar enrichment by 3 different types of stars considered in the chemical evolution model
Figure 16: Cumulative fraction of the contribution of different types of stars to 16O, 20Ne, 32S, 36Ar (the most abundant stable isotopes of each element) enrichment as a function of 12+log(O/H) predicted by the chemical evolution model of Medina-Amayo (2023) and Amayo & Carigi (in preparation). Cyan line: asymptotic giant branch stars (AGBs); dark red line: type Ia supernovae (SNe Ia); magenta line: core collapse supernovae (CCSNe).

### v2

@@ 5.3 Ar/O
Figure 13: log(Ar/O) as a function of 12+log(O/H) (top) and as a function of 12+log(Ar/H) (bottom). Black squares and blue circles represent star-forming galaxies and H ii regions of the DESIRED-E sample, which Ar/O ratios have been calculated using the ICF(Ar) scheme by Izotov et al. (2006). The orange diamonds indicate the mean log(Ar/O) values considering bins in 12+log(O/H) or 12+log(Ar/H). The red continuous lines represent linear fits to the data represented in each panel. The green continuous line shows the time evolution of Ar and O abundances of the ISM predicted by the CEM of the Milky Way of Medina-Amayo (2023). The black dashed lines and the grey bands show the solar 12+log(O/H), 12+log(Ar/H) and log(Ar/O) and their associated uncertainties, respectively, from Asplund et al. (2021).
In Sect. 4.3, we concluded that the different ICF(Ar) schemes considered show a very similar behaviour of the Ar/O ratio with respect to the O abundance and degree of ionisation, so the analysis of the Ar/O ratio seems to be independent of the final selection of the ICF(Ar) scheme. In fact, Fig 14 indicates that the differences in the Ar/H computed using any pair of ICF(Ar) schemes are lower than 0.04 dex when O2+/O ¡ 0.95 but not much larger than 0.10 for higher ionised objects. Due to that, we finally decided to use the ICF(Ar) by Izotov et al. (2006) in order to be consistent with the results obtained with the rest of alpha-elements and to remove the objects with lower ionisation degree (O2+/O $<$ 0.2), where the contribution of unseen Ar+ is larger. This allows the elimination of objects where the ICF(Ar) value is higher and minimizes the systematic uncertainty introduced by the use of an ICF.
@@
In Fig. 13 we represent the log(Ar/O) vs. 12+log(O/H) (top) and vs. 12+log(Ar/H) (bottom) distributions of SFGs+H ii regions of the DESIRED-E sample. We have represented the data points with the same symbol regardless of whether [Ar iii] or [Ar iii]$+$[Ar iv] lines were used to determine the Ar abundance. As in some previous figures, in Fig. 13 we also include a least-squares fit to the data represented in each panel. The two distributions show weak correlations with slopes of different signs. Although the sample Pearson correlation coefficients of the linear fits to log(Ar/O) vs. 12+log(O/H) and vs. 12+log(Ar/H) distributions are r = --0.28 and 0.17, respectively, their very low $p$-values indicate that both correlations can be considered real. As in the case of the fits separated by type of object, the fit to log(Ar/O) vs. 12+log(O/H) for SFGs+H ii regions indicates a decrease of log(Ar/O) of 0.24 dex in the whole interval of 12+log(O/H) represented in the top panel of Fig. 13. This behaviour is contradictory to what is expected from current nucleosynthesis and chemical evolution models. In fact, the curve predicted by the CEM of Medina-Amayo (2023) included in Fig. 13 would imply an increase of the Ar/O ratio with metallicity, in clear disagreement with the general behaviour shown by our data. This puzzling result deserves further study
@@
In the case of the log(Ar/O) vs. 12+log(Ar/H) distribution shown in Fig. 13, we find a weak linear correlation with a small positive slope and a total increase of log(Ar/O) of 0.12 dex in the entire Ar/H range. Although the predictions of the CEM are now qualitatively consistent with the data, they are not quantitatively consistent. The CEM predicts an increase of 0.54 dex over the range of Ar abundances represented in Fig. 13, considerably higher than observed. On the other hand, unlike what happens in the cases of Ne and S, the CEM also does not reproduce the mean log(Ar/O) at low metallicities. The predicted value is about 0.10 dex lower than observed, but this cannot possibly be due to depletion of O onto dust grains because it is expected to be negligible at such low metallicities.
@@
Regarding the log(Ar/O) vs. 12+log(Ar/H) distribution obtained by other authors, the results are disparate. Arellano-Córdova et al. (2024) and Miranda-Pérez & Hidalgo-Gámez (2023) obtain a quite flat distribution in the log(Ar/O) vs. 12+log(Ar/H) relation at intermediate values of 12+log(Ar/H) but opposite trends in low and high values of the Ar abundance. In particular, Miranda-Pérez & Hidalgo-Gámez (2023) find a rather strong increase of log(Ar/O) for objects with 12+log(Ar/H) $>$ 6.0, reaching values larger than in any SFG of our DESIRED-E sample. On the other hand, Arnaboldi et al. (2022) studied the behaviour of the Ar/O vs. Ar/H relation for a sample of planetary nebulae in the M31 galaxy. They find an increase of the Ar/O ratio as the Ar/H increases but in a rather narrow interval of 12+log(Ar/H), between 6.0 and 6.5 approximately, in agreement with the predictions of the chemical evolution models of the Milky Way thick disc and solar neighbourhood of Kobayashi et al. (2020). Arnaboldi et al. (2022) interpret the increase in the log(Ar/O) vs. 12+log(Ar/H) plane they observe as due to the contribution of SNe Ia. In fact, the models by Kobayashi et al. (2020) predict that about 34% of the total solar Ar is produced by SNe Ia. Fig. 16 also indicates that in the CEM of Medina-Amayo (2023) the increase of Ar/O with metallicity is produced basically by SNe Ia. We do not see an Ar/O increase as large as that reported by Arnaboldi et al. (2022), and our results are broadly consistent with a constant log(Ar/O) distribution in most of the 12+log(Ar/H) range. However, the bins represented in the bottom panel of our Fig. 13 suggest a slight increase –of only $\sim$0.1 dex– for values of 12+log(Ar/H) $\gtrsim$ 6.0 which –as we demonstrate in Fig. 12– does not appear to be an artifact of the ICF(Ar). In principle, our linear fit to the log(Ar/O) vs. 12+log(Ar/H) distribution could be interpreted as an indication of a certain contribution by SNe Ia, although in a more limited proportion than that proposed by Arnaboldi et al. (2022).
@@ 6 Conclusions
We derive Ne/O ratios for 1228 DESIRED-E spectra, of those 557 correspond to H ii regions and 671 to SFGs. The log(Ne/O) vs. 12+log(O/H) distribution of H ii regions shows a large dispersion and no evidence of correlation. Our analysis indicates that the position of those objects cannot be correctly reproduced by any of the ICF(Ne) schemes considered. Therefore, we do not recommend the use of these objects for the study of the cosmic evolution of Ne. On the other hand, the SFGs show much lower dispersion in their Ne/O ratios and rather similar log(Ne/O) vs. 12+log(O/H) and log(Ne/O) vs. 12+log(Ne/H) linear relations with positive but small slopes. These correlations may be due to several factors or a combination of them. Firstly, the effect of a metallicity-dependent O depletion fraction that increases throughout the metallicity range covered by the objects. Secondly, an effect of the ICF(Ne), which does not completely cancel out the Ne/O dependence on O2+/O. A third possibility would be a certain unknown mechanism producing Ne, but chemical evolution models do not predict such scenario, at least in a significant amount.
In order to achieve more confident determinations, we only consider objects for which their S/H ratio can be determined from the sum of S+ and S2+ abundances. Considering this, the log(S/O) vs. 12+log(O/H) relationship that we found separately for 441 and 492 spectra of H ii regions and SFG of our sample is compatible with a constant value, especially for H ii regions and when we combine all the sample objects (H ii regions + SFGs), in contrast with some previous works (Dors et al., 2016; Díaz & Zamora, 2022; Brazzini et al., 2024). On this regard, we have found that at least part of the results of Díaz & Zamora (2022) are incorrect, so some of their conclusions should be revised. We find a rather tight and positive linear fit for all the objects when we use the log(S/O) vs. 12+log(S/H) distribution. It is remarkable that the predictions of the CEM by Kobayashi et al. (2020) fit fairly well that distribution in the whole S/H range. We propose that the positive slope the log(S/O) vs. 12+log(S/H) relation and its change of slope at 12+log(S/H) $\sim$ 6.0 can be interpreted as an observational evidence of contribution of S produced by SNe Ia.
@@ Acknowledgements.
Data Availability
The complete tables of atomic data, references, physical conditions and abundances for all the objects in our sample can be found at: https://zenodo.org/records/15005362.
@@ Appendix B The relation between the oxygen abundance and ionisation degree
Figure 15: 12+log(O/H) as a function of the ionisation degree, O2+/O, for DESIRED-E sample of Galactic and extragalactic H ii regions (blue circles, left panel) and star-forming galaxies (black squares, right panel). The red diamonds and lines indicate the mean 12+log(O/H) values and its standard deviation, considering bins of 0.2 in O2+/O.
Appendix C Relative contribution to O, Ne, and Ar enrichment by 3 different types of stars considered in the chemical evolution model
Figure 16: Cumulative fraction of the contribution of different types of stars to 16O, 20Ne, and 36Ar (the most abundant stable isotopes of each element) enrichment as a function of 12+log(O/H) predicted by the chemical evolution model of Medina-Amayo (2023). Cyan line: asymptotic giant branch stars (AGBs); dark red line: type Ia supernovae (SNe Ia); magenta line: core collapse supernovae (CCSNe).